Pfizer (US company) sells a drug in Canada.

  • Sales: CAD 45.0m
  • Variable Costs: USD 10.0m, Euro 1.5m and CAD 2.0m
  • Current exchange rate: CAD 1.25/USD and USD 1.1/Euro.
  1. Calculate the gross margin in USD: ______
  2. Calculate the gross margin in CAD: ______
Answer #1

rate positively ..

Given that-
sales = CAD 45
Variable cost
USD 10
EURO 1.5
CAD 2
USD to CAD exchange rate = 1/1.25 0.80
Euro to CAD exchange rate = 1.1*1.25 1.3750
Ans a) Calculate the gross margin in USD: ______
Local currency Local currency Amount Exchange rate USD AMT
i sales = CAD 45          0.80 36
Variable cost
USD 10          1.00 10
EURO 1.5          1.10 1.65
CAD 2          0.80          1.60
ii Total USD cost = 13.25
iii=i-ii Gross margin in USD = 22.75
Ans b) Calculate the gross margin in CAD…..
Local currency Local currency Amount Exchange rate USD AMT
i sales = CAD 45          1.00 45
Variable cost
USD 10          1.25 12.5
EURO 1.5          1.38 2.0625
CAD 2          1.00          2.00
ii Total USD cost = 16.5625
iii=i-ii Gross margin in CAD = 28.4375